# A Bidirectional Long Short Term Memory Approach for Infrastructure Health Monitoring Using On-board Vibration Response Authors: Reza Riahi Samani

### v1

@@ 1 Introduction
Infrastructure health monitoring is a fundamental component of maintaining a safe and reliable infrastructure network. Continuous monitoring over time and across the infrastructure network facilitates the infrastructure health conditions estimation, which can be used for condition-based maintenance strategies [1, 2]. This process involves assessing infrastructure damage, and tracking changes in physical properties, such as modal frequencies, modal mass, modal damping, stiffness, and mode shapes [3, 4].
@@
Traditionally, infrastructure physical parameter estimation methods relied on sophisticated identification algorithms to solve inverse problems in modal parameter estimation [9, 10, 11]. For example, [10, 11] investigated track stiffness estimation using optimization techniques namely, cross-entropy [10] and adaptive regularization [11] to solve the identification optimization problem and to infer the railway track stiffness profile. However, implementing such algorithms is often computationally expensive and time-consuming [12]. Furthermore, the centralized nature of parametric damage detection methods makes them infeasible for real-time damage detection applications, especially as infrastructure data continues to grow exponentially. Hence, recent advances in machine learning, particularly deep learning techniques, offer promising alternatives for overcoming these challenges [13, 14, 15, 16].
In the current paper, we further explore the estimation of infrastructure physical parameters through deep learning approaches. We consider railway infrastructure to demonstrate the application of our proposed methodology in estimating railway track stiffness parameters using vibration response signals. Railway inspection often relies on Axle Box Acceleration (ABA) monitoring systems[7, 17, 18], hence, we utilize ABA measurements to estimate track stiffness. Furthermore, our methodology holds promise for broader applications in infrastructure inspections utilizing drive-by vibration measurements. For further information on drive-by infrastructure inspections, readers can refer to [19].
@@ 1.1 Deep learning in structure health monitoring using vibration response
Deep learning techniques have increasingly gained attention in infrastructure health monitoring in recent years. The strength of these techniques to detect and locate damage directly from vibration response without any need for data preprocessing or hand-crafted feature extraction is highly promising compared to traditional vibration-based structure monitoring techniques [12]. Deep learning techniques generally require a two-stage process including a damage-sensitive feature extraction through the raw acceleration signals, and processing the extracted features to assess the health state of the structure [12, 20]. Accordingly, an autoencoder-based framework has been proposed by [20] for structural damage identification, comprising two main components: dimensionality reduction and relationship learning. The first component reduces the dimensionality of the original input vector, and the second component learns the relationship between the features and the stiffness reduction. Furthermore, [21] proposed a convolutional neural network (CNN) framework to predict the track dynamic stiffness using the ABA measurements in real-time, including a comparison of the performance between the standard CNN and dilated CNN algorithms. Both models performed well considering their accuracy, with the dilated CNN requiring less computation time in both the training and deployment processes. While CNN effectively captures relevant information within a neighborhood of samples, it often struggles to learn long-term dependencies in sequential datasets, which is relevant for railway track parameter estimation over a long period of data. To address this, [22] proposed a 1DCNN-LSTM-ResNet architecture to identify structural damages based on time-dependent data. The model employs one-dimensional CNN (1DCNN) for feature extraction, Long Short-term Memory (LSTM) for recognizing long-term dependencies, and ResNet to counteract the vanishing gradient problem during deep network training. The proposed architecture outperformed 1DCNN, LSTM, and their combination in diagnosing the damage states of the Z24 bridge located in the Bern district near Solothurn, Switzerland.
@@ 3.1 Problem statement
The proposed approach involves a multi-level sequential data analysis. First, the lower-level network performs the feature extraction task over the vibration response segments. Second, the upper-level BiLSTM network takes the extracted features to estimate the health conditions over the sequence of beams. In the following, Section 3.2 discusses the lower-level bearing span feature extraction methods, followed by Section 3.3, which discusses the upper-level BiLSTM networks for the estimation of health conditions.
@@ 3.2 Lower level bearing span feature extractor
The idea behind using 1DCNN in the feature extraction phase is to leverage the strength of 1DCNN in capturing local relationships in this phase. The 1DCC can provide features for the upper-level model based on local relations within the input signal. The 1DCNN architecture used in this paper comprises three sets of convolutional and max-pooling layers, followed by a flattening layer and a fully connected layer that constructs the output layer of the feature extraction. The hyperparameters of the network include the number of convolutional layers, the number of filters in each layer, filter lengths, and the size of the dense layer, which is the latent space size of the feature extractor. Features extracted from the network are fed as the input for the subsequent model in the upper-level analysis.
@@
3.2.2 LSTM Cell
In contrast to the previous approach, this feature extractor aims to emphasise the temporal relationships within the input sequence. While the vibration signal exhibits long-term temporal relations in the upper-level beam-to-beam context, the LSTM-based feature extractor can also be incorporated to emphasise temporal relations within the feature extraction phase. The feature extractor used in this paper consists of an LSTM layer that takes the frames of vibration signals and returns the hidden state of the last time step as the extracted features for each frame. The extracted features are fed as the input for the subsequent BiLSTM networks in the upper-level analysis. The hyperparameters of the feature extractor include the number of LSTM layers and the number of LSTM units per layer.
@@ 3.3 Upper-level BiLSTM Networks
After extracting features from the bearing span frames in the lower-level analysis, the output is fed into BiLSTM networks at the upper level to capture the long-term temporal relations between sequences of beams. This approach is based on the physics of vehicle-based vibration response signals, where sequences are related to both the forward and backward paths of the sequence. For example, when there is a defect in one of the beams in a sequence, it is reflected not only in the forward path of the drive-by vibration signal but also in the backward path. Therefore, we propose using BiLSTM networks, which consist of stacking two BiLSTM layers, and the outputs of the layers are concatenated and passed to the subsequent layers. The bidirectional information is fed to the final fully connected networks to produce the output vector. The hyperparameters include the number of BiLSTM layers, the units of each layer, and the units of the fully connected networks. The schematic view of the BiLSTM networks is provided in Figure 2.
@@ 4 Illustrative case study of Railway track monitoring
Railway track stiffness is a key indicator of the track’s health and is commonly used to monitor the condition of the infrastructure. This stiffness is primarily influenced by the stiffness of the ballast, but it is also affected by the stiffness of the fastening systems, which include bolts, clamps, and rail pads. In this paper, track stiffness is considered by railpad and ballast stiffness parameters, denoted as the vector $k=[k_{\mathrm{p}},k_{\mathrm{b}}]^{\intercal}$, where the influence of fastening components is considered via railpad stiffness. The objective of the analysis is to estimate the track stiffness using observed Axle-Box Acceleration (ABA) vibration responses.
@@ 4.1 Data
Data preparation for this case study is conducted using the finite element model proposed in [7]. This model accurately represents the physics of the vehicle-track interaction and simulates ABA measurements at an operational speed of 65 km/h by solving the vehicle-track motion equation. The dataset includes cases of constant track stiffness and local changes in track stiffness related to the railpad and ballast stiffness. Constant track stiffness represents the normal condition of railway tracks, where there is no change in track stiffness along the track. Ballast and railpad stiffness values in the simulation model are randomly selected from the stiffness range set R1 in Table 1, using a uniform distribution.
The stiffness reduction is considered in two scenarios: reduction in one sleeper and reduction in three sleepers within a 10-sleeper segment at random locations. The stiffness reduction in one sleeper simulates a defect such as a hanging sleeper, where an unsupported sleeper hangs within the railway segment, causing a local reduction in track stiffness at that sleeper. The stiffness reduction in three sleepers simulates a scenario where a reduction in substructure support over a wider length leads to reduced stiffness in multiple sleepers [7]. The stiffness of the railpad and ballast are simultaneously reduced to the stiffness range set R2 in Table 1, using a random variable with uniform distribution. The stiffness range set R2 represents multiple faults with varying severities, including degraded or missing railpads or local reductions in ballast stiffness, indicating potential component degradation [7]. The random locations of stiffness reduction within the track segment prevent biases in the spatial distribution of defects, making the simulations more representative of real-world scenarios.
@@
To test the robustness of the proposed method in a noisy environment, which is relevant to drive-by vibration response observations, an additive white Gaussian noise $\sim N(0,\sigma^{2})$ is considered for the ABA signal, where $\sigma$ is the standard deviation of the measurement noise. A noise-to-signal ratio of 15% is used, with a variance equal to 15% of the ABA signal power. The additive 15% Gaussian noise in the ABA signal is considered to simulate real-world conditions and evaluate the performance of the proposed model under such circumstances. Each of the noise-free and noise-added scenarios consists of 15 000 records of 10-sleeper track segments, and the datasets are split into training, validation, and test sets, with 60%, 20%, and 20% of the total records, respectively. Figure 3 illustrates the schematic view of a 10 sleeper track segment and the corresponding framing for the vibration response signal.
Figure 3: The layout of the 10 sleepers track segment.
@@ 4.2 Results
During the training and validation process, the data is split into training and validation sets. Figure 4 shows the training process of the four models (CNN-LSTM, LSTM-LSTM, CNN-BiLSTM, and LSTM-BiLSTM) on the noise-added datasets, demonstrating the loss convergence versus the number of epochs on a logarithmic scale. The figure indicates that the LSTM-BiLSTM architecture achieves the lowest training loss compared to the other models. The validation set is used to fine-tune the hyperparameters of the models using a grid search method. The performance of the selected fine-tuned models is then evaluated on the test set, with the results presented in Table 2. Two evaluation metrics, Mean Absolute Percentage Error (MAPE) and Root Mean Square Error (RMSE), are used to compare the accuracy of the models in stiffness estimation. MAPE measures the average absolute percentage difference between predicted and actual values, while RMSE measures the square root of the average squared differences between predicted and actual values.
Figure 4: Training loss convergence in the scenario of the noisy dataset.
@@ 4.3 Discussion
The employed CNN and LSTM feature extractors process vibration response data in the time domain and effectively handle noisy conditions. The proposed method does not require any data preprocessing steps or handcrafted feature extraction based on expert knowledge, providing an automatic health condition monitoring algorithm capable of learning features through the training process. However, since drive-by vibration response signals are non-stationary and their components vary in both time and frequency, transformation to the frequency domain may result in a better representation of the signal. This could improve model performance but would impose further computational costs in preprocessing and may require expert knowledge in the feature extraction phase [30]. Hence, Future research could explore the trade-off between different feature extraction approaches, and enhance the proposed method with the right feature extraction approach.
@@
The illustrative case study considers a railway track infrastructure comprising beams and beam nodes. In this context, the rails and the sleepers function as the beams of the structure, while the locations of the sleepers are considered beam nodes that transfer the load to the ballast substructure. The proposed approach involves segmenting the vibration signal into bearing spans corresponding to individual sleepers or beam nodes. This methodology shows potential for application to other infrastructures with a similar configuration of beams and beam nodes, such as bridges. Furthermore, the dataset used in this paper is based on the vehicle-track interaction finite element simulation [7], enhanced by incorporating floating defect locations and additive white Gaussian noise. Therefore, future research could explore the performance of our proposed LSTM-BiLSTM network in other case studies of infrastructure health monitoring using drive-by vibration response measurements.

### v2

@@ 1 Introduction
Infrastructure health monitoring is a fundamental component of maintaining a safe and reliable infrastructure network. Continuous monitoring over time and across the infrastructure network facilitates the infrastructure health conditions estimation, which can be used for condition-based maintenance strategies [1, 2]. This process involves assessing infrastructure damage and tracking changes in physical properties, such as modal frequencies, modal mass, modal damping, stiffness, and mode shapes [3, 4, 5, 6].
@@
Traditionally, infrastructure physical parameter estimation methods relied on sophisticated identification algorithms to solve inverse problems in modal parameter estimation [11, 12, 13]. For example, [12, 13] investigated track stiffness estimation using optimization techniques namely, cross-entropy [12] and adaptive regularization [13] to solve the identification optimization problem and to infer the railway track stiffness profile. However, implementing such algorithms is often computationally expensive and time-consuming [14]. Furthermore, the centralized nature of parametric damage detection methods makes them infeasible for real-time damage detection applications, especially as infrastructure data continues to grow exponentially. Hence, recent advances in machine learning, particularly deep learning techniques, offer promising alternatives for overcoming these challenges [15, 16, 17, 18, 19, 20].
In the current paper, we further explore the estimation of infrastructure physical parameters through deep learning approaches. Railway inspection often relies on Axle Box Acceleration (ABA) monitoring systems [9, 21, 22]. Hence, we develop a novel deep learning framework to estimate infrastructure health parameters using ABA measurements and demonstrate the application of our model utilizing simulated ABA signals for different track stiffness parameters. Furthermore, our methodology holds promise for broader applications in infrastructure inspections utilizing drive-by vibration measurements. For further information on drive-by infrastructure inspections, readers can refer to [23].
@@ 1.1 Deep learning in structure health monitoring using vibration response
Deep learning techniques have increasingly gained attention in infrastructure health monitoring in recent years. The strength of these techniques to detect and locate damage directly from vibration response without any need for data preprocessing or hand-crafted feature extraction is highly promising compared to traditional vibration-based structure monitoring techniques [14]. Deep learning techniques generally require a two-stage process including a damage-sensitive feature extraction through the raw acceleration signals, and processing the extracted features to assess the health state of the structure [14, 24]. Accordingly, an autoencoder-based framework has been proposed by [24] for structural damage identification, comprising two main components: dimensionality reduction and relationship learning. The first component reduces the dimensionality of the original input vector, and the second component learns the relationship between the features and the stiffness reduction. Furthermore, [25] proposed a convolutional neural network (CNN) framework to predict the track dynamic stiffness using the ABA measurements in real-time, including a comparison of the performance between the standard CNN and dilated CNN algorithms. Both models performed well considering their accuracy, with the dilated CNN requiring less computation time in both the training and deployment processes. While CNN effectively captures relevant information within a neighborhood of samples, it often struggles to learn long-term dependencies in sequential datasets, which is relevant for railway track parameter estimation over a long period of data. To address this, [26] proposed a 1DCNN-LSTM-ResNet architecture to identify structural damages based on time-dependent data. The model employs one-dimensional CNN (1DCNN) for feature extraction, Long Short-term Memory (LSTM) for recognizing long-term dependencies, and ResNet to counteract the vanishing gradient problem during deep network training. The proposed architecture outperformed 1DCNN, LSTM, and their combination in diagnosing the damage states of the Z24 bridge located in the Bern district near Solothurn, Switzerland. In [27], the frequency spectrum of simulated acceleration signals was utilized to develop and train a 1DCNN deep learning algorithm. The paper [27] showed that a one-dimensional variant of VGG convolutional networks can detect bridge damage across a range of real-world noise signals. In fact, convolutional neural networks (CNNs) are designed to efficiently extract spatial and hierarchical features from data, making them particularly effective for inputs like vibration response signals. The process of applying filters and pooling in a CNN helps capture essential features while reducing spatial dimensions, which not only enhances computational efficiency but also preserves critical information [27].
@@ 3.1 Problem statement
The proposed approach involves a multi-level sequential data analysis. First, the lower-level network performs the feature extraction task over the vibration response segments. Second, the upper-level BiLSTM network takes the extracted features to estimate the health conditions over the sequence of beams. The two steps are performed in sequence with an end-to-end approach. This eliminates defining handcrafted features, which may require expert knowledge and limits a fully automated analysis. The hidden layers of the deep learning network are learned through the backpropagation process corresponding to the last layer of the network. The initial layers may extract more meaningful features, like the sharp changes in the input signal, and the deeper layers may extract more abstract and high-level features. Section 3.2 below discusses the lower-level bearing span feature extraction methods, followed by Section 3.3, which discusses the upper-level BiLSTM networks for the estimation of health conditions.
@@ 3.2 Lower level bearing span feature extractor
The idea behind using 1DCNN in the feature extraction phase is to leverage the strength of 1DCNN in capturing local relationships in this phase. The 1DCC can provide features for the upper-level model based on local relations within the input signal. The 1D-CNN architecture used in this paper consists of convolutional and max-pooling layers, followed by a flattening layer and a fully connected layer, which forms the output layer of the feature extraction process. The hyperparameters of the feature extraction network include the number of convolutional layers, the number of filters per layer, filter lengths, and the size of the dense layer, which represents the latent space of the feature extractor.
To optimize the model architecture, grid search is employed for hyperparameter tuning. The resulting network processes high-frequency time-domain signals of length 8,660 and applies three convolutional and max-pooling layers of 32, 64, and 128 filters with a filter length of 8. The output of convolutional layers is activated by a ReLU function to increase the nonlinear expression of networks. In max-pooling layers, the pool size is $1\times 2$ and the stride is two. The CNNs output is flattened and passed through a dense layer with 128 units to generate the final feature vector for the subsequent health condition estimator.
@@
3.2.2 LSTM
In contrast to the previous approach, The LSTM feature extractor captures temporal relationships within the input sequence. While the vibration signal exhibits long-term temporal relations in the upper-level beam-to-beam context, temporal relations can be captured during the feature extraction phase using an LSTM feature extraction layer. The feature extraction network consists of a single LSTM layer, which processes frames of vibration signals and outputs the hidden state of the last time step as the extracted feature for each frame. This hidden state serves as the feature vector for the subsequent BiLSTM health condition estimator networks. The hyperparameters of the LSTM feature extractor networks include the number of LSTM layers and LSTM units per layer. Grid search tuning identified an optimal architecture with one LSTM layer containing 128 LSTM units.
@@ 3.3 Upper-level BiLSTM Networks
After extracting features from the bearing span frames in the lower-level analysis, the output is fed into BiLSTM networks at the upper level to capture the long-term temporal relations between sequences of beams. This approach is based on the physics of vehicle-based vibration response signals, where sequences are related to both the forward and backward paths of the sequence. For example, when there is a defect in one of the beams in a sequence, it is reflected not only in the forward path of the drive-by vibration signal but also in the backward path. Therefore, we propose using BiLSTM networks consisting of two stacked BiLSTM layers with 128 and 64 units, respectively. The outputs of these layers are concatenated and passed to the subsequent layers. The bidirectional information is then fed into a fully connected network with two dense layers consisting of 64 and 2 units, respectively, to produce the final output vector. The key hyperparameters include the number of BiLSTM layers, the units in each layer, and the number of dense layers and the units in each layer. A schematic view of the LSTM-BiLSTM model architecture is provided in Figure 2.
@@ 4 Illustrative case study of Railway track monitoring
Railway track stiffness is a key indicator of track health and is widely used to monitor infrastructure conditions. It is influenced by various parameters associated with track components, including the rail, sleepers, fastening systems, ballast, and track substructure. Among these, ballast stiffness primarily determines overall track stiffness; however, fastening systems—comprising bolts, clamps, and rail pads—also play a significant role. In this study, track stiffness is characterized using railpad and ballast stiffness parameters, represented as the vector $k=[k_{\mathrm{p}},k_{\mathrm{b}}]^{\intercal}$. The influence of fastening components, such as bolts and clamps, is considered through the rail pad stiffness parameter. The primary objective of this analysis is to estimate track stiffness parameters based on observed Axle-Box Acceleration (ABA) vibration responses.
@@ 4.1 Data
Data used in the illustrative case study is based on the finite-element simulation model proposed in [9] and is available online [31]. In this model, the rail and sleepers are meshed using Timoshenko beam elements, while the ballast and railpads are represented as discrete spring-damper pairs. Clamps and bolts are not explicitly modeled; instead, their stiffness is incorporated into the railpad stiffness, a widely accepted simplification in railway track modeling [32, 9, 33]. The wheel is simplified as a rigid mass, and the wheel-rail contact is modeled using a Hertzian spring. In real-world railway networks, different track types exist due to variations in rail types, fastening systems, sleeper types, and sleeper spacing, which the current simulation may be limited to account for [9]. Figure 3 illustrates the schematic view of the finite-element model for a 10-sleeper track segment alongside the corresponding framing approach for the vibration response signal. We use this model to simulate ABA measurements at an operational speed of 65 km/h by considering floating stiffness reduction and additive white Gaussian noise scenarios.
The dataset includes cases of both constant track stiffness and local variations in track stiffness, due to changes in railpad and ballast stiffness. The constant track stiffness scenario represents normal railway track conditions, where stiffness remains uniform along the track. To simulate this scenario, ballast and railpad stiffness values are randomly sampled from the stiffness range set R1 in Table 1, following a uniform distribution. The stiffness reduction is considered in two scenarios: reduction in one sleeper and reduction in three sleepers within a 10-sleeper segment at random locations. The stiffness reduction in one sleeper simulates a defect such as a hanging sleeper, where an unsupported sleeper hangs within the railway segment, causing a local reduction in track stiffness at that sleeper. The stiffness reduction in three sleepers simulates a scenario where a reduction in substructure support over a wider length leads to reduced stiffness in multiple sleepers [9]. The stiffness of the railpad and ballast are simultaneously reduced to the stiffness range set R2 in Table 1, using a random variable with uniform distribution. The stiffness range set R2 represents multiple faults with varying severities, including degraded or missing railroads or local reductions in ballast stiffness [9]. The lower bound of the range set R2 indicates the more harsh scenarios of the component degradation such as ballast crushing or hanging sleepers [34]. The random locations of stiffness reduction within the track segment prevent biases in the spatial distribution of defects, making the simulations more representative of real-world scenarios.
@@
Noise is an inherent aspect of signal measurement, present even in data acquired from highly accurate sensors. Drive-by vibration responses are also affected by measurement noise due to environmental and vehicle-induced noise. Therefore, an additive white Gaussian noise $\sim N(0,\sigma^{2})$ is added to the ABA signal, where $\sigma$ is the standard deviation of the measurement noise. A noise-to-signal ratio of 15% is used, with a variance equal to 15% of the ABA signal power. The additive Gaussian noise in the ABA signal is considered to simulate real-world measurement conditions and evaluate the performance of the proposed model under such conditions [35, 36]. Each of the noise-free and noise-added scenarios consists of 15 000 records of 10-sleeper track segments, and the datasets are split into training, validation, and test sets, with 60%, 20%, and 20% of the total records, respectively.
Figure 3: The layout of the 10 sleepers track segment, adapted from [9].
@@ 4.2 Results
During the training and validation process, the data is split into training and validation sets. Figure 4 shows the training and validation loss convergence of the four models (CNN-LSTM, LSTM-LSTM, CNN-BiLSTM, and LSTM-BiLSTM) on the noise-added datasets, demonstrating the loss convergence versus the number of epochs on a logarithmic scale. The figure indicates that the LSTM-BiLSTM architecture achieves the lowest training and validation loss compared to the other models. Furthermore, the figure shows no evidence of overfitting in the training process. The validation set is used to fine-tune the hyperparameters of the models using a grid search method. The performance of the selected fine-tuned models is then evaluated on the test set, with the results presented in Table 2. Two evaluation metrics, Mean Absolute Percentage Error (MAPE) and Root Mean Square Error (RMSE), are used to compare the accuracy of the models in stiffness estimation. MAPE measures the average absolute percentage difference between predicted and actual values, while RMSE measures the square root of the average squared differences between predicted and actual values.
Figure 4: Training and validation loss convergence for the noisy dataset.
@@ 4.3 Discussion
The employed CNN and LSTM feature extractors process vibration response data in the time domain. In contrast to previous studies that utilize time-frequency transformations of signals as input for the models [27, 24], the proposed method eliminates the need for data preprocessing or handcrafted feature extraction. Instead, it provides an automatic health condition monitoring framework that learns relevant features directly through the training process. In this approach, we rely on the neural network’s model parameters to extract features from raw vibration signals. However, since drive-by vibration response signals are non-stationary, with components varying in both time and frequency, frequency decompositions may provide valuable insights over the signal [41]. Therefore, future research could investigate novel methodologies that integrate spectral information into deep learning frameworks.
The results indicate that the estimation accuracy of all models decreases in the presence of noise. However, the proposed LSTM-BiLSTM model demonstrates greater robustness to noise compared to the other models, achieving the best performance in the noisy dataset. Despite this, its performance still declines when noise is introduced. Our models solely rely on the neural network parameters learning process to deal with the noise. However, recently, preprocessing techniques have been developed that employ denoising techniques to effectively deal with noisy conditions [42]. Therefore, future research could explore techniques to further enhance the model’s resilience to noise and to develop more robust deep-learning frameworks for noisy environments. This is especially important for real-world applications, where noisy data is a common challenge [42]. Moreover, the CNNs employed in the current paper stem from the Vanilla CNNs architecture employed in [27]. This can be compared with other machine learning models and recent CNN architectures like GoogLeNet and ResNet architectures [43].
@@
The illustrative case study considers a railway track infrastructure comprising beams and beam nodes. In this context, the rails and the sleepers function as the beams of the structure, while the locations of the sleepers are considered beam nodes that transfer the load to the ballast substructure. The proposed approach involves segmenting the vibration signal into bearing spans corresponding to individual sleepers or beam nodes. This methodology shows potential for application to other infrastructures with a similar configuration of beams and beam nodes, such as bridges.
The dataset used in this study is derived from the vehicle-track interaction finite-element simulation proposed in [9]. While this case study demonstrates the proposed framework, simulated vehicle-track interaction signals may not fully capture all environmental and operational factors influencing drive-by measurements. For instance, the dataset in the current work assumes a constant measurement speed as the operational condition, whereas real-world drive-by measurements may involve speed variations. Additionally, the simulated dataset represents a single type of railway track structure, whereas real-world railway networks may exhibit structural variations. Therefore, this study opens new research directions for evaluating the performance of the proposed LSTM-BiLSTM network using field measurements of drive-by vibration responses.